J small decided to buy and sell goods from his garage after he lost his job at the end of April 2019. During may 2019 Small purchased goods to the value of R54 000, one third of which was sold for R38 000 during May 2019. Distribution and administration expenses amounted to R3000 and electricity amounted to R800. Show workings to calculate the profit for the month of May 2019
1
Expert's answer
2020-10-23T07:27:51-0400

J Small’s Profit and Loss Account for the Month of May 2019

Rupees          Rupees

Revenue                                                                                               38,000

Costs and Expenses

Cost of goods sold (note1)                                       18,000

Distribution and administration expenses                 3,000

Electricity                                                                    800

Total costs and expenses                                                           21,800 

Profit                                                                                                    16,200


Notes

  1. Cost of goods sold = 1/3 of purchased goods = 1/3 of R54,000 = R18,000